### v4

# A Comparison of Precinct and District Voting Data Using Persistent Homology to Identify Gerrymandering in North Carolina

Abstract
Gerrymandering is one of the biggest threats to American democracy. By manipulating district lines, politicians effectively choose their voters, rather than the other way around. Current gerrymandering identification methods (namely the Polsby-Poppers and Reock scores) focus on the compactness of congressional districts, making them extremely sensitive to physical geography.
To address this gap, we extend Feng & Porter’s 2021 paper, which used the level-set method to turn geographic shapefiles into filtered simplicial complexes, in order to compare precinct level voting data to district level voting data. As precincts are regarded as too small to be gerrymandered, we are able to identify discrepancies between precinct and district level voting data to quantify gerrymandering in the United States.
By comparing the persistent homologies of Democratic voting regions at the precinct and district levels, we detect when areas have been ‘cracked’ (split across multiple districts) or ‘packed’ (compressed into one district) for partisan gain.
This analysis was conducted for North Carolina House of Representatives elections (2012-2024). NC has been redistricted 4 times in the past 10 years, unusually frequent as most states redistrict decennially, making it a valuable case study.
By comparing persistence barcodes at the precinct and district levels (using the bottleneck distance), shows that precinct-level voting patterns do not significantly fluctuate biannually, while district level patterns do, suggesting that shifts are likely a result of redistricting rather than voter behavior, providing strong evidence of gerrymandering.
This research presents a novel application of topological data analysis in evaluating gerrymandering, and shows persistent homology can be useful in discerning gerrymandered districts.
Keywords: gerrymandering, persistent homology, topological data analysis, precinct-level voting data
1  Introduction
1.1  Gerrymandering
Gerrymandering, or the manipulation of political districts for partisan gain, undermines political representation, and is a divisive issue in today’s political climate. The nature of the United States congressional districts allows state legislatures, rather than congressional legislatures, to determine state and district-level boundaries. As the United States has two polarizing political parties, it creates opportunities for political gerrymandering, inhibiting a fair election process.
A mathematical solution to political gerrymandering could help create a more vibrant and fair democracy. “Elections are a way to hold politicians accountable for what their constituency wants” [5] and with democracy being the foundation of American ideology, our government should fully represent the will of the people. Without any settled solution to political gerrymandering, our democracy continues to be unrepresentative.
Current methods to identify gerrymandering are mainly geometric techniques. Widely accepted gerrymandering identification techniques, such as the Reock and Polsby-Popper scores, are very sensitive to physical geography, as a feature of their formula. Ruiz states that "a measure that should be depended on less in terms of gerrymandering is any measure of compactness" [18].
Nevertheless, these metrics are used regularly to determine how ‘gerrymandered’ a district is. Their frequent use as a gerrymandering identification metric sets a precedent for subpar measures to validate the most fundamental aspect of our democracy.
Geometric applications for identifying gerrymandering have been established for over fifty years, yet the use of discrete geometric methods is still relatively new. Consequently, existing analyses utilizing discrete geometry have not been focused on recent election data, with the most recent being an examination of the 2016 election, by using precinct and district level voting data [10]. Furthermore, there are no comparative studies that analyze the persistence diagrams of precincts and districts alongside commonly used gerrymandering identification techniques.
Below, we introduce rationale on why we use precinct-level voting data to identify gerrymandering.
A precinct is a specific area or district that is used for government purposes, for example voting, law enforcement, or the court system. Precincts are often used to divide larger areas into smaller, more manageable units. For example, during an election, voters are assigned to a specific precinct based on where they live. This helps to ensure that each precinct has a manageable number of voters and that everyone is able to vote. In larger states, precinct-level voting data is collected and released to the public. If we follow the assumption that precincts are too small to be gerrymandered, analysis of precincts will give us insights into how voting patterns relate to district and state-level polling.
The comparison of precinct-level voting data and district-level voting data can be accomplished using persistent homology, a mathematical technique used to detect topological features over a varying scale parameter.
Currently there is little evidence to support or refute the effectiveness of persistent homology as a method for identifying gerrymandering in congressional districts. My research aims to address this gap, and assess the validity of using such a metric to identify gerrymandering.
1.2  Overview
The paper proceeds as follows. We introduce some intuition pertaining to precinct and district level voting data to lay the groundwork for the rest of the paper. In section 2 we introduce relevant background information and introduce Topological Data Analysis (TDA) as well as Persistent Homology (PH). Subsection 2.2, introduces methods for the construction of simplicial complexes, as well as understanding how they relate to identifying gerrymandering. Subsection 2.3 highlights why TDA is best suited to perform this analysis, as well as the novelty of the project. Section 4 introduces specific methods used in this analysis, as well the pipeline for work. Section 5 holistically reviews NC gerrymandering and provides some rationale for our results. Section 7 discusses future directions for the computation of PH as it relates to gerrymandering, and 9 includes generated images from our analysis, which is in North Carolina precincts and districts from 2012-2024.
2  Related Work
Duchin [6] explores the use of generating multiple ensemble maps by using Markov Chains, and comparing the output to actual districts. This offers a gerrymandering identification technique that is not solely based on compactness scores, but has yet to be implemented alongside current redistricting criteria.
2.1  Topological Data Analysis
Topology Data Analysis (and as we will soon see, Persistent Homology) aims to "take data which might be in some higher dimension, and allows us to find the topological features in such a way that we can characterize" [8]. This higher dimensional data comes in the form of point clouds, which, as you might expect, is a collection of data points in 3D space (or any $n$-dimensional space) that might be represented as an object or a shape. Some might be easy to identify, like this point cloud of a torus:
Figure 1: Point Cloud of a Torus
But as we go into higher and higher dimensions, it becomes very hard to visualize the point cloud in 3D. To calculate the persistent homology, we will need to use a filtration of simplicial complexes to analyze our point cloud at various scales of resolution. There are many filtrations possible. One takes each point of the point cloud and makes it a sphere with a diameter $\epsilon$. This diameter is then increased to a certain point, and certain topological features are observed at different scales.
Figure 2: A image of a point cloud filtration
The most intuitive way to understand persistent homology more generally might be to look at research of mapping the brain using Topological Data Analysis (TDA). Dr. Henry Markam at the Blue Brain Project represented neurons as ’vertices’ and neuronal connections as ’edges’ [14]. This relationship between vertices and edges is called a simplicial complex, and can span any dimension. A 0-simplex is just a point, a 1-simplex is a line, a 2-simplex is a filled triangle, and a 3-simplex is a solid tetrahedron. In general, an abstract simplicial complex $K$ is a collection of non-empty subsets of $V_{n}]=\{v_{o},v_{1},v_{3},...,v_{n}\}$ with $n$ as a nonnegative integer $\geq 0$ such that
  1. 1.
    if $\sigma\in K$ and $\tau\subseteq\sigma$, then $\tau\in K$
  2. 2.
    $\{v_{i}\}\in K$ for every $v_{i}\in[v_{n}]$.
Or, for any number $k$ the $k$ simplex is made out of $k+1$ vertices, is k1k-1 dimensional. Algebraic topology allows us to manipulate these simplicial complexes, and observe in higher dimensional spaces.
A 0-simplex is represented by a vertex.
Figure 3: A 0-simplex
Figure 4: A $1$-simplex
Figure 5: A $2$-simplex
Homology, in the topological sense, allows us to determine the number of ’holes’ in data [19]. Persistent homology allows us to interpret how the holes in the data change over time, when a certain scale parameter $\epsilon$ is changed.
Homology aims to count the number of holes in a simplicial complex - whether it is a two-dimensional hole:
ABC
Figure 6: A the boundary of a 22-simplex
Or a three-dimensional void:
ABCD
Figure 7: A 33 dimensional simplicial complex
To calculate the homology of a simplicial complex, we use the Betti numbers, which indicate how many holes of each dimension an object has. The Betti number is a topological invariant, which means it does not change under continuous transformations. Because these numbers are independent of the topology of the space (the exact position of points may not matter, only their relative positions in relation to each other), so, we can study the topology of a space without explicit geometric description. Take the torus:
Figure 8: Image of a torus
The first Betti number counts the number of one dimensional, circular, holes. which in the case of the torus would be two. The next Betti number represents the amount of two dimensional holes, which in the case of the Torus is one (the hollow void inside of the Torus is the 2nd dimensional hole).
The Betti numbers can also be applied to simplical complexes.
  • $\beta_{0}$: Number of connected components.
  • $\beta_{1}$: Number of independent loops or 1-dimensional holes.
  • $\beta_{2}$: Number of independent voids or 2-dimensional holes
The $k$-th Betti number $\beta_{k}$ counts the number of $k$-dimensional cycles that are not boundaries.
Consider the following simplical complex:
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$
Figure 9: 11-dimensional simplicial complex
  • $\beta_{0}=2$: There are 2 connected components (two separate simplical complexes).
  • $\beta_{1}=8$: There are 8 1D holes (all edges form closed simplices).
  • $\beta_{2}=2$: There are 0 2D holes (hollow interiors).
Persistence barcodes arise in the field of topological data analysis (TDA. While Betti numbers summarize the topological features of a simplicial complex (the number of connected components, holes, etc,), persistence barcodes extend this idea to study how these features evolve across different scales in a dataset.
Barcodes are a visual representation of persistent homology, which tracks the birth and death of topological features as we move through a filtration of simplicial complexes.
2.1.1 Persistent Homology
Persistence barcodes extend the concept of Betti numbers by studying how features persist or disappear as we vary the scale of the dataset. This is particularly useful in analyzing point clouds or filtered simplicial complexes.
  • Filtration: Imagine starting with 0-simplicies (just vertices) and gradually adding edges, triangles, and higher-dimensional simplices as we increase a scale parameter $\epsilon$.
  • At each step, the topology changes:
    • New connected components may form ($\beta_{0}$).
    • 1D holes may form and then be filled ($\beta_{1}$).
    • 2D holes may appear and then close ($\beta_{2}$).
2.1.2 Persistence Barcodes
A barcode is a collection of intervals, one for each topological feature, showing when the feature appears (birth) and when it disappears (death) during the filtration. Recall that the Betti numbers are topological invariants. This idea is particularly useful when applied to persistence barcodes, as the persistence of topological features in the barcode reflects meaningful topological properties of the space or data set, independent from the specific geometry. This allows for studying the shape of data without being concerned about specific coordinates.
Persistent features (those that appear at multiple scales) are typically more reliable, as they are topologically significant.
Non-persistent features appear at only a small scale and vanish quickly.
  • Long bars represent persistent features (likely impor tant structure).
  • Short bars often represent noise or small-scale fluctuations.
  1. 1.
    We start with a Point Cloud: Consider a set of points in $\mathbb{R}^{2}$. Initially, each point is its own connected component ($\beta_{0}=$ number of points).
  2. 2.
    Build a Filtration: Gradually increase $\epsilon$, adding edges when the $\epsilon$ areas intersect, then triangles, and higher simplices:
    • At small $\epsilon$: Points connect into clusters (changes in $\beta_{0}$).
    • At medium $\epsilon$: Loops form as points connect but don’t yet fill in triangles ($\beta_{1}>0$).
    • At large $\epsilon$: Loops fill in, and connected components merge.
  3. 3.
    Compute Persistent Homology: For each $k$-dimensional feature ($k=0,1,2$), track when it appears (birth) and disappears (death).
  4. 4.
    Visualize as Barcodes:
    • $\beta_{0}$ barcode: Tracks the persistence of connected components.
    • $\beta_{1}$ barcode: Tracks the persistence of 1D holes.
    • $\beta_{2}$ barcode: Tracks the persistence of two dimensional holes
This can be applied to any $n$ dimensions
The next question that arises is how this $\epsilon$ affects the persistence barcode. We will visualize how a simplicial complex evolves under a filtration process. As the scale $\epsilon$ increases, new simplicial complexes are added, and topological features emerge.
The filtration process is typically visualized by showing a sequence of simplicial complexes at different scales. At each step, we observe how the features of the complex evolve as $\epsilon$ increases.
For each change in $\epsilon_{i}$, the homology of the filtered simplicial complex is calculated. The homology $(H_{0})$ for the amount of connected components, $(H_{1})$ for the amount of $1$-dimensional holes, etc. The birth and death of features is also described in persistent homology. These are specific features that appear and disappear when we analyze the data across different filtrations. A birth is when new features (e.g., a connected component) first appears in the data. For example, if our filtration was a Vietoris-Rips filtration, (increasing the size of circles around points in a point cloud, edge is created between two points when the circle intersect) the first time a certain edge $e$ is formed between points $a$ and $b$ would be the birth of $e$ [8]. Conversely, the death of a feature is when it ceases to exist. As we keep on increasing our scale $\epsilon$, loops or connected components may merge, causing the original features to disappear. This would mark the death of the features.
2.2  Review Of Literature
2.2.1 Topological Data Analysis (TDA)
Persistent homology is the most widely recognized tool used in the burgeoning field of topological data analysis (TDA), where concepts from algebraic topology are used to simplify, summarize and compare complex datasets. Methods from topological data analysis have proven successful in neuroscience [3], [4], medical diagnostics [16], and machine learning [1] among other applications. However, applications to demographic or electoral geospatial data seem to be very limited.
2.2.2 Topological Data Analysis as it relates to gerrymandering
Feng and Porter [8] proposed a method to implement the use of persistent homology in precincts. Using released precinct voting data from the 2016 cycle, they generated maps using this data. The darker the intensity of the color, the more votes for each candidate (Clinton and Trump).
In order to analyze the simplicial complex over a ‘time,’ or a certain parameter, it goes through a filtering process, and then the persistence is taken from it. The construction is as follows:
Let $X=X_{0}\subset X_{1}\subset\cdots\subset X_{l}$ be a filtered simplicial complex. The $m$th persistent homology of $X$ is the pair
$\left(\{H_{m}(K_{i})\}_{1\leq i\leq l},\{f_{i,j}\}_{1\leq i\leq j\leq l}\right),$
where $f_{i,j}:H_{m}(X_{i})\to H_{m}(X_{j})$, for all $i\leq j$ and $m$ smaller than the dimension of $X$, are the maps that are induced by the action of the homology functor on the inclusion maps $X_{i}\hookrightarrow X_{j}$. We refer to the collection of all $m$th persistent homologies as the persistent homology (PH) of $X$ [8].
The challenge arises when it comes to how the filtered simplicial complex will be constructed. As demonstarted by Feng and Porter, widely used constructions for filtered simplicial complexes are not adequate for visual-level data. “Traditional distance-based constructions on the LA Times voting data yield ambiguous results about the persistence of features" [8]. Specifically, the Vietoris–Rips complexes and Alpha complexes slightly increase the distance in each filtration of a point-cloud data set. As voting-level data and a point-cloud data set have very little resemblance, these constructions do not allow for the preservation of distance in precinct-level maps. We are looking to compare precinct-level maps and voting data with district-level maps and voting data, meaning that preserving real distance between precincts is vital. Feng and Porter propose two novel techniques to construct filtered simplicial complexes: queen adjacency and rook adjacency.“Two precincts are queen-adjacent if they touch at any two points, including corners, distinct from ‘rook adjacency,’ in which two precincts are adjacent if they share a boundary" [8].
The filtering is defined as:
$\delta_{b,r}(p)=\frac{|V_{b}(p)-V_{r}(p)|}{V_{b}(p)+V_{r}(p)},$
where $V_{b}(p)$ is the number of blue (i.e., Biden) votes in a precinct $p$ and $V_{r}(p)$ is the number of red (i.e., Trump) votes in that precinct. For the first network, they consider only those precincts for which $\delta_{b,r}(p)\geq 0.95$, for example. Each time, this value is reduced. For the next network in the sequence, we take all precincts with $\delta_{b,r}(p)\geq 0.90$. We continue until they consider all precincts in which Clinton won.
In the original creation of the precinct-level map, precincts were colored either red or blue to indicate the voting data in that precinct. However, in order to analyze the voting data over time, only one group of data (either Trump or Clinton) was filtered. In this case, Democratic votes were not colored and were analyzed as ‘holes’ in the data, with the use of $H_{0}$ and $H_{1}$, the 0th and 11st Betti numbers. The second method was the Level-Set Method, which will be used in this paper. The Feng and Porter paper concluded that both methods for filtering simplicial complexes were better than conventionally used filtering methods for geospatial voting data.
Using the adjacency complex, the filtration was the margin of Republican victory in a precinct or district. Say, starting with red precincts with a 90% win margin, and then lower the threshold (and more precincts would be added). Then, you would get homological features where the stronger the persistence of the feature, the larger the difference between within the precinct, and the surrounding areas. This allows us to detect voting islands, and in this paper, see where these islands have been manipulated.
These methods were applied to precinct-level voting data in comparison with district-level voting data in order to identify gerrymandering. This general idea is demonstrated in an expository paper by Kallal [10]. Kallal states that if a district is not gerrymandered, it should accurately represent precinct-level voting data at the district level. In persistent homology, this would be represented as “the district-level barcode being similar to the precinct-level barcode." A barcode in persistent homology encodes information about the birth and death of topological features present in a filtration of simplicial complexes as $\epsilon$ changes. Kallal introduces analyses pertaining to comparisons of precinct and district-level barcodes.
2.3  Novelty
As well as this extension to Kallal’s work, I intend to do a novel comparison of the results from precinct - district level voting analysis to other conventionally used gerrymandering techniques, in order to isolate the strengths and weaknesses district measure.
Duchin explains how to incorporate topological data analysis in Markov Chain Monte Carlo technique [7]. Duchin visualizes and analyzes large ensembles of computer-generated distracting plans to understand gerrymandering in enacted maps. Duchin emphasizes that persistent homology is not primarily useful for classifying plans, while reinforcing the narrative of robustly identifying relationships between geography and aggregated vote data. Hence, in my research, I am focusing on actual enacted plans with actual voting data to draw conclusions.
Kallal focuses on presidential elections data for the sake of consistency. However, district-level diagrams that he generates do not necessarily match with the actual results of the House of Representative elections. As an example, when I classify 2024 North Carolina Presidential and Congressional election data, President Trump won majority 11/13 congressional districts; however, Democrats won 4 congressional seats. Specifically, North Carolina Congressional District 1 was won by President Trump by a 17,793 vote margin (202,762 votes vs. 184,969 votes). However, in the congressional elections, the Democratic candidate won the district, presumably because of the Libertarian party candidate.
Figure 10: North Carolina District 1 2024 Results
To avoid similar split-ticket issues, I solely focus my research on congressional races. There is no publicly available source for congressional elections (like the MGGG-States database that Duchin and Kallal focuses on), so I decided to create my own dataset, involving creating precinct and district level maps.
In the state of North Carolina, county boards are required to provide vote tabulation data by precincts, allowing me to conduct analysis of precincts. There is currently no analysis being done at the precinct (and district) level using TDA. Additionally, no analysis spans the breadth of data as this paper does, as I collect and analyze data from 2012 - 2024, looking at the biannual US House Election. This gives us novel insight into how gerrymandering is carried out in one of the most gerrymandered states (North Carolina), and how persistent homology stacks up against other widely used gerrymandering identification techniques.
3  Objective
We aim, as Feng and Porter did, to find the phenomenon of a ‘blue island in a sea of red’ [8]. In elections, cities tend to more blue (Democratic voting) and surrounded by this ‘sea,’ or large, sparsely populated Republican voting areas. We treat these voting island as ‘holes’ in the data, and can therefore compute the homology for these features. Once we detect these voting islands, we can see if and when they have been altered in relation to their surrounding environment, providing a clearer picture when it comes to understanding gerrymandering.
4  Methods
The methodology of the paper is as follows.
4.1  Data Collection
To conduct analysis, I downloaded historical congressional US House of Representatives shapefiles from North Carolina State Board of Elections (https://www.ncsbe.gov/). A shapefile is a vector data format that stores geographic features’ location, shape, and attributes. Shapefiles are a common file format for geospatial data and are often used in geographic information systems (GIS) software.
North Carolina Election Board publishes a csv file post election by precinct. A sample CSV file extract is provided below. The file is very comprehensive and shows county precinct, election date, contest type, candidates, party, election day, absentee, provisional and total votes.
Figure 11: North Carolina 2024 Precinct Voting Data
We use an R code to filter these data into table as below. We use “US House of Representatives District” Contest name and sum votes by Republican and Democratic candidates by precincts. For the purpose of our analysis, we discard third-party results.
Figure 12: North Carolina 2024 Precinct Voting Data (in R)
Once, we have data by precinct, we overlay the results in the Precinct shapefile in R. This data is then verified using QGIS, then exported as a shapefile for data analysis.
When doing this analysis, we encountered certain missing and/or corrupted data. North Carolina has $2,716$ precincts and in certain elections we had data from between 1-50 precincts missing (on average about 0.01 precincts didn’t have data). For simplicity we just fill this data with 10 votes for Democrats and 10 votes for Republicans and continue with our analysis.
Figure 13: Pipeline of data collection and processing
We undertake similar work for districts. Fortunately, Wikipedia maintains a comprehensive list of past elections, congressional districts, and votes by candidate in each of the election results. We take each elections shapefile and manually overlay this voting data in QGIS to create the district shapefile.
4.2  Data Analysis
Next, this data was converted to in a raster format (.tiff), to which we can apply Kallal’s modification to get a filtered simplicial complex.
Data analysis will be conducted through Python code designed for analyzing political precinct and district data by constructing a level set complex using a modified level set algorithm based on the work of Feng and Porter, and updated by Kallal. It follows the below pipeline:
shapefile $\rightarrow$ some raster images $\rightarrow$ a filtered simplicial complex $\rightarrow$ a barcode
Figure 14: Diagram showing data processing pipeline
4.2.1 Workflow
Margin and Density Calculation: The Republican and Democratic margins are created by the code, and scaled based on the area of whichever precinct they come from, mitigating errors that arise. If we were to skip this step, “the construction of the simplicial complex would not [be] sensitive to variations in the absolute voter margin per square meter […] Specifically, a region with only a small margin one way or the other gets filled in by the expanding levelset at the same rate as a region of the same size," [10] which would not yield an accurate barcode. The actual computing of the margins are done using GDAL [9] and OGR libraries, which allow for analysis of GIS (Geographic information system) data embedded in Shapefiles.
Rasterization: We find which precinct/districts are won by Democrats/Republicans, and that specific area is colored blue or red respectively. The result is exported as a .tiff file.
Figure 15: North Carolina 2024 Precinct map (colored)
Level Set Construction: The Level-set construction is done by reducing the win margins (the $\delta$,) of the Democratic vote share. The reason being is that Republicans vote shares generally span large urban areas, where smaller cities are generally more democratic leaning. Therefore, in persistent homology ’holes’ are represented as Democratic leaning precincts. The change in the margins allows us to see which precincts are the most persistent, as they will stand out the most regardless of the change in margins. In this case, more persistent precincts represent precincts that were won by a larger Democratic vote margin.
Next, we are interested in the zero level set, where $\phi=0$. This is essentially the boundary where political affiliation changes (like from Republican to Democrat). Points on this boundary are called vertices. By identifying these points, we can track them over each step of the level set’s evolution. This helps us capture the shape and location of political boundaries at each stage.
These vertices need to be connected to represent the spatial relationships between adjacent points on the boundary. The adjacency list keeps track of which vertices are next to each other, allowing us to create a connected network (a kind of “roadmap” of the boundary points). Once the vertices and their relationships are identified, we use them to build a simplicial complex, a mathematical structure that models these relationships in multiple dimensions (e.g., 1D edges, 2D faces). Now that we have the simplicial complex, we can analyze how it changes over the margin ($\epsilon$). Using the PHAT [2] library, we calculate persistence pairs, which represent the “birth” and “death” of features (clusters) as the level set evolves. These pairs show us which clusters are temporary and which are stable, and which ones are strong outliers.
Figure 16: NC 2024 Precinct Raster 1
Figure 17: NC 2024 Precinct Raster 4
Figure 18: NC 2024 Precinct Raster 9
Figure 19: NC 2024 Precinct Raster 13
Using PHAT to make the persistence barcode
Figure 20: Example of vertices (used to generate barcode)
These vertices are generated by PHAT (Persistent Homology Algorithm Toolkit), and then plotted in a $H_{1}$ (11st Betti number) barcode using matplot.
Figure 21: Example Barcode generated using GUDHI
Bottleneck Distance
The bottleneck distance between two persistence barcodes $B_{1}$ and $B_{2}$ measures the largest possible difference between pairs of points, where each point represents the birth and death of a feature in a topological space. The smaller is distance, the ’closer’ the two barcode are to one another. Let $\gamma:B_{1}\rightarrow B_{2}$ be a bijection (pairing) between the points in $B_{1}$ and $B_{2}$. The bottleneck distance is defined as:
$d_{B}(B_{1},B_{2})=\inf_{\gamma}\max_{x\in B_{1}}\|x-\gamma(x)\|_{\infty}$
where $\|x-\gamma(x)\|_{\infty}$ denotes the $L_{\infty}$-norm (or maximum distance) between matched points $x\in B_{1}$ and $\gamma(x)\in B_{2}$. This metric is particularly useful for capturing the largest single difference between barcodes. This was calculated for precinct and district barcode for each biannual House Election (2012, 2014, etc.), using GUDHI.[13]
Below are results for precinct to district bottleneck distance, as well as precinct to precinct (biannaually), bottleneck distance.
Figure 22: Precinct to District Bottleneck Distance (2012-2014)
Figure 23: Precinct to District Bottleneck Distance
Here we see little precinct-precinct fluctation biannually, signifying that there is not a massive demographic change. However, precinct - district change shows that change is with redistricting, rather than voting shift.
Comparision to other metrics: The Polsby-Poppers and Reock scores were calculated for North Carolina’s 2022 map, and its 2024 map. A two-tailed, paired t-test was used, and a statistically significant result was found between the two years. This validates our analysis that 2022 and 2024 maps had a significant change in their redistricting.
Figure 24: Polsby-Poppers and Reock Scores (2022 and 2024)
Figure 25: P-Value for two-tailed, paired, T-Test
Total Persistence:
The total persistence of a barcode measures the aggregate lifespan of features within a single barcode rather than directly comparing two barcodes; it compares the sum of the lengths of all the births and deaths in a barcode. For a given barcode $B$ with points $(b_{i},d_{i})$, [11] the total persistence $\text{TP}(B)$ with respect to the $p$-norm is given by:
$\text{Total Persistence}(B)=\sum_{i}(d_{i}-b_{i})^{p}$
In order for $TP(B)$ to be finite, it is necessary to replace any infinite values with a finite number, such as the maximum possible death time [11]. This was calculated for all barcodes at all timepoints.
5  Analysis and Discussion
5.1  History of North Carolina Gerrymandering:
North Carolina’s gerrymandering goes back to early 1990s, when the state legislature created the highly irregular and predominantly African American 12th congressional district in the 1992 redistricting process. Legal challenges surrounding this district ultimately led to U.S. Supreme Court case Shaw v. Reno in 1993 where in 5-4 decision, the U.S. Supreme Court ruled that the 12th congressional district violated the Equal Protection Clause of the Fourteenth Amendment.  [17] In more recent history, 2011 route of statewide redistricting occurred following the release of official 2010 census counts in March of 2011, and new plans were enacted for 2012 and 2014 elections.
Congressional districts were redrawn in February 2016 after a challenge from United States District Court for the Middle District of North Carolina.
More recently, in October 2023, North Carolina lawmakers proposed two separate plans for redistricting, plans 756 and plans 757 which resulted in considerable media attention and outrage. In fact, Prof. Mattingly commented on his blog that newly proposed North Carolina maps are more gerrymandered and less responsive than maps struck down in 2021 [15]. This is most recently seen in the 2024 election results, where "North Carolina Democrats won over 46 percent of the vote in congressional races (after adjusting for uncontested races), but a mere 29 percent of congressional seats." [12]
Summarizing, the North Carolina Precincts have remained unchanged while the US House of Representatives plans were changed four times since the 2011 elections, resulting in a very interesting real-life test case for this research.
  • 2024 Map – Enacted in 2023
  • 2022 Map – Enacted in 2021 post 2020 Census (added one seat)
  • 2020, 2018, 2016 – Same maps, enacted in 2016 post court challenge
  • 2014, 2012 – same maps, enacted in 2011.
5.2  Rational
Precincts are the most accurate reflection of how areas vote, purely based on their small size. If we can compare the persistence diagrams of district, and the comparative precincts, it would give us a better understanding of how and when districts have been manipulated. Persistent homology offers the perspective of understanding if votes are accurate representations when it comes to evaluating them on a district level. ‘If the state is not gerrymandered, we should expect the district level barcode to be similar to the precinct level barcode’ [10], and a lower Bottleneck distance. If there is a certain precinct that is very persistent, but that same persistence is not represented in the district level barcode, there is reason to believe that the district has been manipulated, and cracking may be at play. On the other hand, if the district level barcode is showing persistence that is not reflected in the precinct level barcode, mapmakers might be employing packing in order to consolidating votes in fewer districts.
From 2012 - 2024, the NC precinct boundaries did not change at all. By definition, they should not change unless for a massive demographic change, which is not seen in North Carolina during this time period. This lack of change at the local precinct level provides a stable base for analyzing voting patterns. However, while the precinct boundaries themselves stayed the same, there were broader shifts in political trends that aligned with national or state-wide political movements, which caused fluctuations in voting behavior. As expected, the bottleneck distance between precinct level barcodes did not change significantly.
Yet, redistricting at the district level displays a significant level of change. In North Carolina, like in many other states, redistricting occurs every ten years following the census, but in practice, district boundaries can be adjusted more frequently, often in response to shifts in population, by court action, or as a result of partisan gerrymandering.
The difference between precinct-level constancy and the district-level fluctuation shows that redistricting in North Carolina does not always align with precinct-level voting behavior. This raises questions about how effectively district-level redistricting represented the true political makeup of the states.
North Carolina 2012:
The 2012 redistricting cycle in North Carolina, "following majorities in both houses of the General Assembly in the November 2010 elections for Republicans" had noticeable altered district boundaries, and led to a disproportionate win for Republicans (a gain of 3 out of 13 contested seats). By looking at both the precinct and district level, it is evident that the district votes do not fully represent precinct level votes.
Figure 26: NC 2012 Precincts won colored
Figure 27: NC 2012 Precinct Barcode
Figure 28: NC Levelset #1
Figure 29: NC Levelset #25
The one precinct that persists at the end is in the Charlotte metropolitan area, which rationally makes sense. Charlotte is our blue ‘hole’ in the ‘sea’ of red. However, this is not reflected in the district level barcode. The district containing Charlotte (district 12), will die after the 12th iteration of the levelset (out of a total of 25), a clear instance of cracking (diluting) the more democratic leaning voters of Charlotte. Additionally, voters in North Carolina’s 1st and 4th district were at a loss, as both their districts died at levelset 10 out of 25.
Figure 30: NC 2012 Districts won colored
Figure 31: NC 2012 Precinct Barcode
Figure 32: NC District Levelset #1
Figure 33: NC District Levelset #7
North Carolina 2022:
During the 2021 redistricting cycle, the Charlotte metropolitan area was split into 2 democratic districts, both of which were persistent in the district level data. On the precinct level data, as seen in 2012, Charlotte was persistent along with Raleigh, a democratic leaning college-town. Democrats won 7 out of 13, their highest total in over a decade.
The persistence of these two Democratic-leaning districts from the Charlotte metro area allowed Democrats to gain a stronger foothold in North Carolina, which led to more Democratic candidates winning state legislative and congressional seat following the 2022 election. This was a stark contrast to the earlier part of the 2010s when Republicans had held strong control. However, in response to the gains Democrats made in the 2022 elections, Republicans in North Carolina began pushing for new redistricting plans ahead of the 2024 elections. Since North Carolina is a key battleground state, control of the legislature is critical for both parties, as it determines how congressional and state legislative districts are drawn, which leads us to the 2024 election.
Figure 34: NC 2022 Precinct won colored
Figure 35: NC 2022 Precinct Barcode
Figure 36: NC Precinct Levelset #1
Figure 37: NC Precinct Levelset #20
Figure 38: NC 2022 District won colored
Figure 39: NC 2022 District Barcode
Figure 40: NC District Levelset #1
Figure 41: NC District Levelset #20
North Carolina 2024: During the 2024 election, Republicans presented two plans (746 and 747), which were critiqued heavily (by both academics and the media) for being overtly partisan. The enacted plan, 747, is rated by the Princeton Gerrymandering Project as an F [20]. The only persistent district is district 12, which contains Charlotte, and there are no barcodes that persist until the end of the 25 levelset iterations. The fact that the district containing Charlotte is not persistent compared to 2022 (where both districts contained half of Charlotte were both persistent), could be a artifact of the ‘red-wave’ in 2024, or partisan redistricting. Additionally, district 2, containing Raleigh, died during iteration 15 of the level set, whereas the precincts around Raleigh died around levelset 20. This implies that Democrats have been effectively “cracked" into specific districts (to minimize their influence everywhere. If we look to NC 2022, the reduction of one democratic seat should not have led to them losing 3 seats in the House. The phenomenon cannot even be attributed to the nationwide ‘red sweep.’ For example, in the case of North Carolina district 1, Trump carried the state, but Don Davis - a Democrat - retained his seat. The two very persistent districts support the idea that Democrat support has been diluted across other districts, leading Republicans to flip 3 seats in the 2024 election.
Figure 42: NC 2024 Precinct won colored
Figure 43: NC 2024 Precinct Barcode
Figure 44: NC 2024 District won colored
Figure 45: NC 2024 District Barcode
Bottleneck Distance
Table I: Bottleneck Distance, Precinct vs, Precinct, 2012 - 2024
Table II: Bottleneck Distance, District vs, District, 2012 - 2024
We calculate the bottleneck distance between district and precinct barcodes for each year. There are minimal changes in precinct-level barcodes across year; stable voting patterns, but significant changes in district-level barcodes, driven by redistricting processes.
Total Persistence
The total persistence was also calculated, but needs more work to determine its usefulness in identifying gerrymandering.
Table III: Table including total persistence for every 2 years 2012-2024
Comparison to other gerrymandering identification measures
Figure 46: Reock and Polsby-Poppers scores for NC 2022 Districts
Figure 47: Reock and Polsby-Poppers scores for NC 2024 Districts
As seen, the two state-mandated measures for identifying gerrymandering in North Carolina have no significant difference between 2022 and 2024, even though Republicans flipped 3 seats. These measures are highly sensitive to geography, and many times do not detect partisan redistricting, but rather interesting geographic phenomenon. This shows the need for a new measure to be able to identify gerrymandering where others fail.
6  Conclusion
In conclusion, North Carolina’s redistricting process over the years has been a contentious issue, with visible inconsistencies between precinct and district representations suggesting potential gerrymandering. Despite legal and academic scrutiny, district boundaries have been frequently altered, often in response to political pressures, thus not always accurately reflecting local voting behavior. Persistent homology analysis and the bottleneck distances have started revealing discrepancies that traditional methods might miss, emphasizing the urgent need for improved gerrymandering detection techniques.
7  Future Work
The Wasserstein distance is another measure that computes a cumulative distance between pairs of points in two persistence barcodes. For a given $p\geq 1$, the $p$-Wasserstein distance is defined as:
$d_{W,p}(B_{1},B_{2})=\left(\inf_{\gamma}\sum_{x\in B_{1}}\|x-\gamma(x)\|^{p}\right)^{\frac{1}{p}}$
In this case, the distance considers all differences between matched points rather than focusing on the largest single difference. This could be used in future work.
  1. 1
    Duchin uses Fréchet means to compare different maps; this is something that could be implemented.
  2. 2
    We don’t consider population shifts, growths, racial makeup change etc. in drawing congressional boundaries. We strictly rely upon actual enacted maps. We hope to investigate either of these approaches in future work.
8  Acknowledgments
I would like to science research teacher, Ms. Dardis, for her continued support through this research. I would like to thank Dr. Ismar Volić, from Wellesley College for his input of my work, as well as the Institute of Mathematics and Democracy. Thanks to Mr. Kallal for his guidance in generating raster images and understanding the methods of the project. Thanks to Dr. Mason Porter and Dr. Thomas Weighill for their input on my project.
This work is supported by the Carnegie Foundation and Carnegie Young Leaders. Thank you for your support.
All errors are my own, feel free to reach out with corrections, comments, and suggestions at ananya.neytri.shah@gmail.com.
9  Additional Figures
All code is available at https://github.com/an-shah/persistenthomology. Additional figures are below.
Figure 48: NC 2012 Precinct Raster
Figure 49: NC 2012 Precinct Barcode
Figure 50: NC 2012 Precinct Level Set 1
Figure 51: NC 2012 Precinct Level Set 25
Figure 52: NC 2012 District Raster
Figure 53: NC 2012 District Barcode
Figure 54: NC 2012 District Level Set 1
Figure 55: NC 2012 District Level Set 25
Figure 56: NC 2014 Precinct Raster
Figure 57: NC 2014 Precinct Barcode
Figure 58: NC 2014 Precinct Level Set 1
Figure 59: NC 2014 Precinct Level Set 25
Figure 60: NC 2014 District Raster
Figure 61: NC 2014 District Barcode
Figure 62: NC 2014 District Level Set 1
Figure 63: NC 2014 District Level Set 25
Figure 64: NC 2016 Precinct Raster
Figure 65: NC 2016 Precinct Barcode
Figure 66: NC 2016 Precinct Level Set 1
Figure 67: NC 2016 Precinct Level Set 25
Figure 68: NC 2016 District Raster
Figure 69: NC 2016 District Barcode
Figure 70: NC 2016 District Level Set 1
Figure 71: NC 2016 District Level Set 25
Figure 72: NC 2018 Precinct Raster
Figure 73: NC 2018 Precinct Barcode
Figure 74: NC 2018 Precinct Level Set 1
Figure 75: NC 2018 Precinct Level Set 25
Figure 76: NC 2018 District Raster
Figure 77: NC 2018 District Barcode
Figure 78: NC 2018 District Level Set 1
Figure 79: NC 2018 District Level Set 25
Figure 80: NC 2020 Precinct Raster
Figure 81: NC 2020 Precinct Barcode
Figure 82: NC 2020 Precinct Level Set 1
Figure 83: NC 2020 Precinct Level Set 25
Figure 84: NC 2020 District Raster
Figure 85: NC 2020 District Barcode
Figure 86: NC 2020 District Level Set 1
Figure 87: NC 2020 District Level Set 25
Figure 88: NC 2022 Precinct Raster
Figure 89: NC 2022 Precinct Barcode
Figure 90: NC 2022 Precinct Level Set 1
Figure 91: NC 2022 Precinct Level Set 25
Figure 92: NC 2022 District Raster
Figure 93: NC 2022 District Barcode
Figure 94: NC 2022 District Level Set 1
Figure 95: NC 2022 District Level Set 25
Figure 96: NC 2024 Precinct Raster
Figure 97: NC 2024 Precinct Barcode
Figure 98: NC 2024 Precinct Level Set 1
Figure 99: NC 2024 Precinct Level Set 25
Figure 100: NC 2024 District Raster
Figure 101: NC 2024 District Barcode
Figure 102: NC 2024 District Level Set 1
Figure 103: NC 2024 District Level Set 25